OpenStudy (anonymous):

Line AB contains points A (−2, 6) and B (4, 5). The slope of line AB is −6 negative 1 over 6 1 over 6 6

OpenStudy (anonymous):

negative 1 over 6

OpenStudy (anonymous):

thankyou!! :D

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@crabby_patty do you think you could help me with 2 more?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok cool thanks the first one: Point G is located at (3, −1) and point H is located at (−2, 3). Find the x value for the point that is 2 over 3 the distance from point G to point H. 1.5 −0.33 −0.5 6.33

OpenStudy (anonymous):

sry don't know this one

OpenStudy (anonymous):

its ok:) heres the second one: The equation of line EF is y = 2x + 1. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, 2). y = 2x − 4 y = 2x + 2 y = negative 1 over 2x − 4 y = 1 over 2x + 2

OpenStudy (anonymous):

it is either the first or second one

OpenStudy (anonymous):

lol thats what i thought. thankyou!

OpenStudy (anonymous):

yup